sonya375:

Mrs. Ming invested an amount of money in two accounts for one year. She invested some at 8% interest and the rest at 6% interest. Her total amount invested was $1,500. At the end of the year, she had earned $106.40 in interest. How much had Mrs. Ming invested in the account paying 6%?

Vocaloid:

let s = the amount invested in the 6% account let e = the amount invested in the 8% account "total amount = 1,500" means s + e = 1500 amount earned in interest = (interest rate) * (amount invested) so applying this to both accounts 0.06*s + 0.08*e = 106.40 so putting both our equations together s + e = 1500 0.06*s + 0.08*e = 106.40 solve for s

Vocaloid:

@sonya375 still there? let me know if you're having trouble solving the system